5) Write down a power series expansion (centered at r = 0) for the function f(x)
Answer #1

Solution: Powen senies expansion of ex centered ato is em = 1+ 2+ entre sie mi-x+ - .. fl^)= Je*dx = Jl com ens de 24-Stussy